OpenStudy (anonymous):

NEED AN ANSWER ASAP ! WILL GIVE MEDALL !! The graph of a line has a slope of 2 and passes through the point (–6, –5). What is the equation of the line in point-slope form? y – 6 = 2(x – 5) y + 6 = 2(x + 5) y – 5 = 2(x – 6) y + 5 = 2(x + 6)

OpenStudy (solomonzelman):

y=mx+b m is slope b is the y-intercept substitute 2 for m, y=2x+b find the y-intercept, by plugging in your point. (–6, –5) -5=2(-6)+b -5+12=b b=7 So y=2x+7, or \[B\]

OpenStudy (anonymous):

option b ?